OpenStudy (anonymous):

2x+9=5x-12 can someone please help me??

OpenStudy (skyphos):

This is English...lol

OpenStudy (skyphos):

Anyway... what you do is take away 5x on the right

OpenStudy (skyphos):

so 5x-5x cancels each other out..so your equation should look like, 2x-5x+9=-12 (because what you do on one side, you do to the other)

OpenStudy (skyphos):

you subtract 2x-5x, you get -3x....so -3x+9=-12

OpenStudy (skyphos):

You isolate -3x by getting rid of 9 by subtracting it, so 9-9=0..... then you do the same on the other side...so -3x=-21

OpenStudy (skyphos):

divide by -3 on both sides and x=7

OpenStudy (skyphos):

:D

OpenStudy (anonymous):

next time post on math!

OpenStudy (kohai):

Math questions belong in math chat